OpenStudy (anonymous):

Andrea's body temp. dropped from 103.2 degrees farenheit at 3 am to 99.7 degrees farenheit at 8 am. Find the rate of change in her body temperature.

OpenStudy (anonymous):

(T2-T1)/(t2-t1)=(103.2-99.7)/(8-3)= 3.5/5=0.7

OpenStudy (anonymous):

where T= temperature, t=time

OpenStudy (anonymous):

0.7 fahrenhite per hour